OpenStudy (anonymous):

Consider the equation 7x + 3y = 42. Part 1: On your own paper, graph this equation using the slope-intercept method. In the space provided, explain, in words, each step of the procedure you used. Make sure to use complete sentences and correct grammar. (3 points) Part 2: On your own paper, graph this equation using the intercepts method. In the space provided, explain, in words, each step of the procedure you used. Make sure to use complete sentences and correct grammar.

OpenStudy (anonymous):

(1) 7x+3y=42 3y=-7x+42 y=-7/3+14 m= -7/3 and y-intercept(b)=14 , we can use the point (0,14) to graph, given the (0,14), plot it to the cartesian plane, and given the slope, from (0,14), move 7 units down since it's negative, and 3 units to the right considering that it's positive. .

OpenStudy (stacey):

y=-7/3 x +14